OpenStudy (anonymous):

For f(x)=(x+2)^4(x-3), 2 is a zero of multiplicity 4. True or False. If false why?

OpenStudy (campbell_st):

well to find the zeros you need to solve x + 2 = 0 and x - 3 = 0 multiplicity means that the (x + 2) is raised to the power 4... its like saying (x+2)(x+2)(x+2)(x+2)(x-3) hope this helps

OpenStudy (anonymous):

so then I would have to foil them out. then use synthetic division? i know the answer already, but i was wondering if there were an easier way?

OpenStudy (anonymous):

i just realized the answer is in the question and that means it's -2 right off the back

OpenStudy (campbell_st):

nope... all you need to do is solve x + 2 =0 x - 3 = 0 and then decide is x = 2 is a zero of of multiplicity 4 is you have a quaratic f(x) = (x + 3)(x - 5) the zeros are when x + 3 = 0 which is x = -3 and when x - 5 = 0 which is x = 5 so apply the same reasoning... (x + 2)^4 means its a repeated root.... repeated 4 times...

OpenStudy (anonymous):

sorry thank you!!!

OpenStudy (campbell_st):

correct... so you were asked if 2 is a zero... its false..

OpenStudy (anonymous):

yerp yerp yerp thank you(:

OpenStudy (whpalmer4):

@joshuaknmcguire that should be "yerp yerp yerp yerp" — it's a yerp of multiplicity 4, you know :-)